# Can We Catch the Elephant? The Evolvement of Hallucination Evaluation on Natural Language Generation: A Survey

### v1

@@ Abstract
Hallucination in Natural Language Generation (NLG) is like the elephant in the room, obvious but often overlooked until recent achievements significantly improved the fluency and grammatical accuracy of generated text. For Large Language Models (LLMs), hallucinations can happen in various downstream tasks and casual conversations, which need accurate assessment to enhance reliability and safety. However, current studies on hallucination evaluation vary greatly, and people still find it difficult to sort out and select the most appropriate evaluation methods. Moreover, as NLP research gradually shifts to the domain of LLMs, it brings new challenges to this direction. This paper provides a comprehensive survey on the evolvement of hallucination evaluation methods, aiming to address three key aspects: 1) Diverse definitions and granularity of facts; 2) The categories of automatic evaluators and their applicability; 3) Unresolved issues and future directions.
@@ 1 Introduction
The hallucination problem in Natural Language Generation (NLG) typically refers to situations where the generated text deviates from the facts in source or external knowledge. Like the elephant in the room, this problem has existed since the beginning of the development of NLG. However, in the era of Large Language Models (LLMs), with the rapid improvement of fluency and grammatical accuracy in text generation, this issue has gradually gained more attention. The evaluation of hallucination motivates the development of LLMs to be more reliable and safe. In this paper, we make a comprehensive survey on the hallucination evaluation methods and elaborate them according to the development process, which gives a broad view of current trends and future directions.
@@
The definition of hallucination varies a lot from work to work over time. In the initial stages, hallucination refers to the occurrence of semantically incorrect but grammatically viable content (Lee et al., 2018). This setup is typically rooted in traditional tasks such as Text Summarization and Neural Machine Translation (NMT), which often involve definite source information. The breakthrough came with the advent of LLMs like ChatGPT (OpenAI, 2022), crossing the barriers of conversational systems. These models are capable of processing various text inputs and providing responses. In this case, users have more expectations for the responses, especially the consistency of world facts under open answers. Many previous tasks can also be accomplished within a conversation by prompting LLMs with correct instructions (Ouyang et al., 2022). In the response of LLMs, hallucination occurs now and then, with a mix of faithful and factual problems.
@@
Faithfulness and factuality are two concepts that are closely related when describing hallucinations in the generated text and can be prone to be confusing in some circumstances. In this paper, we add prefixes to both of them for better understanding by introducing Source Faithfulness (SF) and World Factuality (WF)  [^1] In this paper, we also write SF as faithfulness/consistency, and WF as factuality.. SF measures the degree to which the generated output accurately reflects and is consistent with the source input. It ensures that the output stays true to the information provided in the source. SF has a more limited scope, as there are specific source documents that can be used to substantiate and verify the generated text. WF, on the other hand, assesses whether the generated output aligns with general world knowledge and facts. WF is a more expansive and challenging problem as it goes beyond the specific source and considers the broader context of common sense and established knowledge, which are more difficult to comprehensively collect and encode. Recent studies have recognised the critical importance of addressing and improving the SF and WF of generated text.
@@
Whether in existing NLG tasks or the more practical scenarios in the LLMs era, SF and WF problems manifest in different forms. For example, in the NMT task, generated translations detached from the source text are considered factually incorrect (Dale et al., 2023). In the summarization task, generated content usually should belong to the source document to be faithful, but some also argue that certain hallucinations can align with external facts (Dong et al., 2022). In the dialogue task, it is important to balance the hallucination and user experience (Santhanam et al., 2021). In LLM downstream tasks, the issue is often related to the prompt, which is designed for different scenarios and to retrieve the knowledge that existed in LLM. Apart from these, LLM faces unique difficulties, such as updated world information and false-premise questions (Vu et al., 2023; Kasai et al., 2024).
@@
Hallucinations are prevalent across various tasks, as it is not a problem that has been completely defined, yet it intuitively pertains to binary correctness. While humans can readily perceive some hallucinations, the model often struggles to do so. Moreover, in the current development of LLMs regarding reasoning, planning, task adaptability, etc., hallucination evaluation is an issue that requires careful consideration. Previous works have some introductions on benchmarks for LLM hallucinations (Huang et al., 2023; Zhang et al., 2023c; Ji et al., 2023; Huang et al., 2021), but they do not include evaluators before LLMs and more research after, thus resulting in a lack of comparison between developments before and after. Also, they do not specifically focus on evaluators, therefore lacking in-depth analysis of details, as well as their respective strengths and weaknesses.
@@
Our contributions lie in three folds. Firstly, this paper organizes automatic evaluation methods for hallucination in the evolvement timeline, including traditional approaches as well as novel methods employing or assessing LLMs, and conducts analytical comparisons. Secondly, an applicable classification method is proposed, categorizing evaluation methods from a practical perspective based on the scenario. Lastly, recommendations for future hallucination evaluations are provided from multiple perspectives. We also give a brief introduction of related fact-checking and human evaluation works in Appendix A and B.
2 Definition and Taxonomy
To understand how to evaluate hallucination in text generation, we first need to define fact. In previous works, fact granularity is a way to define facts and serves as the foundational basis for assessing hallucination. Fact error classification also can help provide a solid and comprehensive perspective for evaluation.
@@ 2.1 Fact Granularity
The definition of fact granularity varies, lacking an agreed-upon labelling scheme (Honovich et al., 2022). Moreover, in generated text, many contents are difficult to have clear definitions. Among previous works, fact granularity can be categorized into the following types.
Token/Word At the token/word level, entities and nouns are the most frequently used granularity, primarily due to their intuitively informative and relative ease of acquisition and comparison. For example, some summarization models tend to forget entities (Maynez et al., 2020; Nan et al., 2021), making up entities that never appeared in the source document.
Span Spans require more annotation than individual tokens or words. Mostly, they are entity-centric, relation-centric, or event-centric, which are closely related to dependency parsing (Goyal & Durrett, 2020; 2021). In evaluators, spans can be used to generate questions (Durmus et al., 2020; Honovich et al., 2021) or check entailment (Kamoi et al., 2023).
Sentence A higher level is sentence, which is coarsely classified whether it contains hallucination or not (Wang et al., 2020; Pagnoni et al., 2021). Sentence-level facts can help obtain an overall perception of the factual content within the sentence.
Passage Passage-level facts can contribute to an overall assessment of the generated text (Fabbri et al., 2021). This encompasses discourse fact connections (Zhang et al., 2023b) and an estimation of passage themes. Moreover, this is crucial for future assessments of the factual accuracy in generating longer texts.
@@ 2.2 Fact Error Type
The taxonomies of fact errors vary subjecting to scenarios, some categorize them by the information source as intrinsic and extrinsic hallucinations (Maynez et al., 2020; Goyal & Durrett, 2021; Cao & Wang, 2021), For faithful error types,  Goyal & Durrett (2021) designs entity-, event-, and noun phrase-related errors. Frank (Pagnoni et al., 2021) categorizes with a more fine-grained scale, including PredE, EntE, CircE, CorefE, LinkE, OutE and GramE. We summarize that such errors can be categorized into Source Faithful Error and World Factual Error, in which way it can cover more NLG tasks including traditional ones and LLM scenarios. SF and WF errors are independent as shown in Figure 1.
Figure 1: Source Faithful Error and World Factual Error examples.
@@ 3 Evaluation Methods Before LLMs
This section contains the automatic evaluation methods for hallucinations in NLG before LLM, which differ from other evaluators that focus on text similarity or grammar correctness. Specifically, the focus mostly lies on the SF aspect of text generation. While previous works rely on the reference to train the models, the reference itself can suffer a lot of factual errors (Maynez et al., 2020), which hinder the model training from learning faithfulness. Also due to the costly but error-prone reference, many works study reference-free evaluation methods, which only require the input of source documents. To introduce the approaches, we will elaborate from evaluators to benchmarks.
@@
It is noteworthy that a significant portion of these evaluation methods is dedicated to assessing summarization tasks, due to which has a fixed source document as an information reference, making faithfulness relatively straightforward to evaluate. Additionally, summarization is commonly employed in fields such as news, medicine, and law, where reliability holds particular importance. Besides summarization, other tasks, such as text simplification, NMT, and dialogue generation, are also included in this section. Some evaluation methods concentrate on a single task, while others can be extended to similar ones.
3.1 With-reference Evaluators
For the summarization task, the approaches start with statistical methods. $\mathbf{Fact_{acc}}$ (Goodrich et al., 2019) uses information extraction models to measure factual overlap. The facts are pre-defined schema as triplet tuples (entity-relation-entity). Then the score is calculated by comparing the schema between ground-truths and generated summaries. Later, Kryscinski et al. (2020) propose a weakly-supervised model FactCC for verifying factual consistency of generated summaries. FactCC uses NLI learning structure to do fact classification. Additionally, the other version FactCCX can highlight spans in the source document that contain the support or opposition for the claim. Besides utilising entity-level facts, DAE (Goyal & Durrett, 2020) points out the dependency-level factual error, and designs a new entailment model. It views dependency arcs as semantic units and applies a 2-class classification task with automatically generated labels $\in\{entailed,non-entailed\}$, where entailing means the semantic relationship it implies between a dependency arc head and child is entailed by the ground-truth sentence, and vice versa.
Text simplification tasks aim to ”translate” texts to a more readable version without altering their meaning. Maskeval (Liu et al., 2022) proposed an automatic evaluator for summarization and simplification tasks. These two tasks share some nature in common, which is they both have long input and output text that need to be loyal to the source. It gets weights from masked LM for every token and multiplies the weight and the Exact Matching (EM) score to get the final score. The experiments show that weights trained on summarization can improve the metric’s performance on simplification, which again confirms the similarity between the two tasks.
@@ 3.1 With-reference Evaluators
Hallucinations in NMT have also received some attention so far (Lee et al., 2018; Raunak et al., 2021).  Guerreiro et al. (2023) proposes a hallucination detection method using glass-box uncertainty-based detectors. This paper states that hallucination in NMT is so rare that like a needle in a haystack.  Dale et al. (2023) also provide a detection method by computing the percentage of the source contribution to the generated translation. It includes reference-based oracles, and both internal and external methods to measure the hallucination percentage.
@@
The above approaches provide a promising starting, as they explore existing datasets and deliver feasible solutions. For long-input tasks such as summarization and simplification, they can use label information, with manual correction making it more accurate. However, a notable drawback is the requirement for annotation, which is costly and labour-intensive, and may also need cross-checking. Additionally, for NMT, the source text serves as the reference. Distinguishing between normal translation errors and hallucinations presents significant challenges for models as well.
@@ 3.2 Reference-free Evaluators
Reference-free means no direct ground truth are provided while evaluating. For abstractive summarization, Question Generation and Question Answer (QA-QG) is a possible pipeline to deal with it. FEQA (Durmus et al., 2020) is an evaluation framework that first extracts question-answer pairs from the summary and then uses an off-the-shelf QA model to extract answers from the document to check if the answers are consistent. QAGS (Wang et al., 2020) follows the framework of FEQA. It uses a fine-tuned BART on NewsQA (Trischler et al., 2017) for QG, which uses cloze-style questions and uses a fine-tuned BERT on SQuAD2.0 (Rajpurkar et al., 2018) for QA. QuestEval (Scialom et al., 2021) is a QA-QG pipeline as well. What’s different is that it lays out a comprehensive evaluation of four dimensions (consistency, coherence, fluency, and relevance). It borrows the QAGS framework for computing the precision score for the summary from the source document and uses SummaQA (Scialom et al., 2019) to compute the recall score, where a question weighted applies. QAFactEval (Fabbri et al., 2022) analyzes all components of the QA-based metric pipeline before, including answer selection, question generation, question answering, answer overlap evaluation, and question filtering. Among them, QAFactEval picks the best components through sound experiments and has made a more powerful combination. To avoid unanswerable-question noise, it also applies a $IsAnsweredSummFilter$ to score and filter the questions. Furthermore, the experiments show that entailment and QA-based metrics offer complementary signals that the combination can achieve better performance. MQAG (Manakul et al., 2023a) measures the consistency between summary and source document by computing the statistical distance of answers over automatically generated multiple-choice questions. The multiple-choice system allows for a more abstract understanding of information and makes it convenient for standard information-theoretic measures. Counterfactual estimation can also be used to evaluate factual consistency. CoCo (Xie et al., 2021) calculates the score by measuring the causal effects from source documents to generated summaries. It can also be generalized to multi-level granularity and other tasks.
Extractive summarization just extracts sentences or sentence spans from source documents. ExtEval (Zhang et al., 2023b) is the first to reveal that extractive summarization also can not guarantee faithfulness. It defines five types of broad unfaithfulness problems which the entailment model can not cover all. Additionally, ExtEval uses four sub-metrics summation to get the unfaithful score.
@@
For Dialog Generation, $\mathbf{Q^{2}}$ (Honovich et al., 2021) is a evaluation method for neural knowledge-grounded generative models. It incorporates QG, QA, and NLI models to score the model’s output to reflect its consistency with its grounding source knowledge. Specifically, it generates questions from named entities and gets answers from the knowledge. Finally, it measures the answer span similarity using an NLI model.
@@
For multiple tasks, AlignScore (Zha et al., 2023) can give a unified function for 7 tasks consistency evaluation. It breaks the context into chunks and the claim into sentences to calculate the alignment score. WeCheck (Wu et al., 2023) finds previous metrics may fail to generalize to real generated samples due to the gap between training and testing and the scarcity of labelled data. It provides a weak supervision learning paradigm based on previous metrics and uses a mixture NLI datasets to warm up and do a noise-aware fine-tuning to get the target metric model.
@@
In comparison to with-reference methods, these approaches have greater generality. It is noticeable that QG-QA models take a large portion. Because, without references, NLI models require accurate extraction of evidence from the source text, which is challenging. Essentially, the QG-QA framework is a flexible approach that extracts and compares information from generated output and source input. When comparing, some methods also incorporate entailment models for judgment. However, the QG and QA steps may introduce considerable noise, which can interfere with the results.

### v2

# Can We Catch the Elephant? A Survey of the Evolvement of Hallucination Evaluation on Natural Language Generation

@@ Abstract
Hallucination in Natural Language Generation (NLG) is like the elephant in the room, obvious but often overlooked until recent achievements significantly improved the fluency and grammaticality of generated text. As the capabilities of text generation models have improved, researchers have begun to pay more attention to the phenomenon of hallucination. Despite significant progress in this field in recent years, the evaluation system for hallucination is complex and diverse, lacking clear organization. We are the first to comprehensively survey how various evaluation methods have evolved with the development of text generation models from three dimensions, including hallucinated fact granularity, evaluator design principles, and assessment facets. This survey aims to help researchers identify current limitations in hallucination evaluation and highlight future research directions.
[…]
Can We Catch the Elephant? A Survey of the Evolvement of Hallucination Evaluation on Natural Language Generation
Siya Qi, Yulan He, Zheng Yuan King’s College London, London, UK {siya.qi, yulan.he, zheng.yuan}@kcl.ac.uk
@@ 1 Introduction
Hallucination in the context of Natural Language Generation (NLG) typically refers to situations where the generated text is inconsistent with, or unsupported by, the input source or external knowledge. Like an elephant in the room, this problem has existed since the beginning of the development of NLG. As text generation models continue to evolve, technologies like Large Language Models (LLMs) have achieved grammatical correctness and fluency nearly indistinguishable from human writing. Consequently, hallucination has gradually surfaced and attracted increased attention. The evaluation of hallucination matters since it motivates the development of LLMs to be more reliable and safe. In this paper, we make a comprehensive survey on the evolution of hallucination evaluation methods through multiple perspectives as in Figure 1, which gives a broad view of current trends and future directions.
@@
The definition of hallucination varies a lot from work to work over time. In the initial stages, hallucination referred to the occurrence of semantically incorrect but grammatically viable content (Lee et al., 2018). This setup was typically rooted in traditional tasks such as Summarization and Neural Machine Translation (NMT), which often involve definite source information. The breakthrough came with the advent of LLMs like ChatGPT (OpenAI, 2022), crossing the barriers of conversational systems. These models are capable of processing various text inputs and providing responses. In this case, users have more expectations for the responses, especially the consistency of world facts under open answers. Many previous tasks can also be accomplished within a conversation by prompting LLMs with correct instructions (Ouyang et al., 2022). In the responses of LLMs, hallucination occurs now and then, with a mix of faithful and factual problems.
Figure 1: The distribution of evaluators on various aspects, including the timeline, hallucinated fact granularity, method design principles, and assessment facets.
@@
Faithfulness and factuality are two concepts that are closely related when describing hallucinations in generated text and can be prone to confusion in some circumstances. In this paper, we add prefixes to both of them for better understanding by introducing Source Faithfulness (SF) and World Factuality (WF). [^1] In this paper, we also write SF as faithfulness/consistency, and WF as factuality. SF measures the degree to which the generated output accurately reflects and is consistent with the source input. It ensures that the output stays true to the information provided in the source. SF has a limited scope, as there are specific sources that can be used to substantiate and verify the generated text. WF, on the other hand, assesses whether the generated output aligns with general world knowledge and facts. WF is a more expansive and challenging problem as it goes beyond the specific source and considers the broader context of common sense and established knowledge, which is more difficult to collect and encode comprehensively. Recent studies have recognized the critical importance of addressing and improving the SF and WF of generated text.
@@
Assessing from SF or WF aspects means the evaluators refer to different source information, which is closely tied to specific tasks. For example, in NMT, generated translations detached from the source text are considered factually incorrect (Dale et al., 2023). In summarization, summaries usually should be faithful to the source document, but some also argue that certain hallucinations can align with external facts (Dong et al., 2022). In dialogue generation, it is important to balance hallucination and user experience (Santhanam et al., 2021). In LLM downstream tasks, the issue is often related to the prompts, which are designed for different scenarios and to retrieve the knowledge that exists in LLMs. Apart from these, LLMs face unique difficulties, such as updating world information and handling false-premise questions (Vu et al., 2023; Kasai et al., 2024).
@@
Hallucinations are prevalent across various tasks, as the problem has not been completely defined, yet it intuitively pertains to binary correctness. While humans can readily perceive some hallucinations, models often struggle to do so. Moreover, in the current development of LLMs regarding reasoning, planning, task adaptability, etc., hallucination evaluation is an issue that requires careful consideration. Previous works have some introductions on benchmarks for LLM hallucinations (Huang et al., 2023; Zhang et al., 2023c; Ji et al., 2023; Huang et al., 2021), but they do not include evaluators before LLMs and more recent research, thus resulting in a lack of comparison between developments before and after. Also, they do not specifically focus on evaluators, therefore lacking in-depth analysis of details, as well as their strengths and weaknesses.
@@
Our contributions lie in three folds. Firstly, this paper organizes automatic evaluation methods for hallucination [^2] We also give a brief introduction of related fact-checking and human evaluation works in Appendix A and B. in the evolution timeline, including traditional approaches as well as novel methods employing or assessing LLMs, and conducts analytical comparisons. Secondly, this paper reviews various evaluation methods from three perspectives: hallucinated fact granularity, evaluator design principles, and assessment facets, highlighting the intrinsic connections and development trajectories among these methods. Finally, we summarize the current state of research on hallucination evaluation, outlining existing issues and potential future directions.
2 Fact Definition
To evaluate hallucination in text generation, we first need to define what constitutes a fact. Previous works have evaluated hallucination by considering different levels of fact granularities and categorizing fact errors into various types.
@@ 2.1 Fact Granularity
The definition of fact granularity varies, lacking an agreed-upon labeling scheme (Honovich et al., 2022), and many elements in generated text are difficult to define clearly. Previous works categorize fact granularity into the following types.
Figure 2: Source Faithful Error and World Factual Error examples.
Token/Word: At the token/word level, entities and nouns are the most frequently used granularities, primarily due to their intuitively informative nature and relative ease of acquisition and comparison. For example, some summarization models tend to forget (Maynez et al., 2020; Nan et al., 2021) or make up entities that never appeared in the source document.
Span: Spans require more annotation than individual tokens or words. Mostly, they are entity-centric, relation-centric, or event-centric, closely related to dependency parsing (Goyal and Durrett, 2020, 2021). In evaluators, spans can be used to generate questions (Durmus et al., 2020; Honovich et al., 2021) or check entailment (Kamoi et al., 2023).
Sentence: A higher level of granularity is the sentence, which is coarsely classified as containing hallucination or not (Wang et al., 2020; Pagnoni et al., 2021). Sentence-level facts provide an overall perception of the factual content within the sentence, allowing for a more holistic assessment.
Passage: Passage-level facts contribute to an overall assessment of the generated text (Fabbri et al., 2021), encompassing discourse fact connections (Zhang et al., 2023b) and an estimation of passage themes. This level of granularity is crucial for evaluating longer texts.
Figure 1 illustrates that evaluation methods before the advent of LLMs mostly relied on text segments, whereas many methods developed after LLMs frequently conduct direct assessments of the entire input text.
@@ 2.2 Fact Error Type
The taxonomies of fact errors vary depending on the scenarios. Some categorize them by the information source as intrinsic and extrinsic hallucinations (Maynez et al., 2020; Goyal and Durrett, 2021; Cao and Wang, 2021). For faithful error types, Goyal and Durrett (2021) designs entity-, event-, and noun phrase-related errors, while Pagnoni et al. (2021) categorize errors on a more fine-grained scale, including PredE, EntE, CircE, CorefE, LinkE, OutE and GramE. We propose categorizing errors into SF Error and WF Error, which can cover a wider range of NLG task, including both traditional ones and LLM scenarios. SF and WF errors can be assessed independently, as illustrated in Figure 2.
@@ 3 Evaluation Methods Before LLMs
This section discusses automatic evaluation methods for hallucinations in NLG before the era of LLMs. These methods differ from those that focus on text similarity (Lin and Hovy, 2003) or grammatical correctness (Bryant et al., 2017), as their primary focus lies on the SF aspect of text generation. Previous works rely on references to train models, but these references can contain factual errors (Maynez et al., 2020), hindering the model’s ability to learn faithfulness. Due to the costly and error-prone nature of references, many works study reference-free evaluation methods that only require the input of source documents.
@@
A significant portion of these evaluation methods is dedicated to assessing the summarization task, which has a fixed source document as an information reference, making faithfulness relatively straightforward to evaluate. Summarization is commonly employed in fields such as news, medicine, and law, where reliability is of particular importance. In addition to summarization, this section covers tasks such as text simplification, NMT, and dialogue generation. Some evaluation methods concentrate on a single task, while others can be extended to similar ones. In the following subsections, we will introduce the approaches by elaborating on evaluators and benchmarks.
3.1 Reference-based Evaluators
Reference-based evaluators compare generated text with ground-truth references to assess factual consistency. For summarization, early approaches start with statistical methods. $\mathbf{Fact_{acc}}$ (Goodrich et al., 2019) uses information extraction models to measure factual overlap based on pre-defined schemas. FactCC and FactCCX Kryscinski et al. (2020) use Natural Language Inference (NLI) models for fact classification and span highlighting. Besides utilising entity-level facts, DAE (Goyal and Durrett, 2020) focuses on dependency-level factual error using entailment models.
Text simplification aims to “translate” texts to a more readable version without altering their meaning. Maskeval (Liu et al., 2022) evaluates both summarization and simplification by combining masked LM weights with Exact Matching (EM) scores, exploiting the similarity between the two tasks.
@@ 3.1 Reference-based Evaluators
Hallucinations in NMT have received attention as well (Lee et al., 2018; Raunak et al., 2021). Hallucination in NMT is rare, like finding a needle in a haystack. It is important to identify the hallucinations with the errors.  Guerreiro et al. (2023) propose a glass-box uncertainty-based detection method, while  Dale et al. (2023) measure hallucination by computing the percentage of source contribution to the generated translation.
@@
These approaches provide promising solutions, as they explore existing datasets and deliver feasible solutions. For long-input tasks such as summarization and simplification, they can use label information, with manual correction making it more accurate. However, a notable drawback is the requirement for annotation, which is costly and labor-intensive, and may also need cross-checking. Additionally, distinguishing between normal translation errors and hallucinations in NMT presents significant challenges.
@@ 3.2 Reference-free Evaluators
Reference-free evaluation methods assess the quality of generated text without relying on direct ground truth. For abstractive summarization, the Question Generation and Question Answer (QG-QA) pipeline is a common approach. FEQA (Durmus et al., 2020), QAGS (Wang et al., 2020), and QuestEval (Scialom et al., 2021) follow a similar framework: extracting question-answer pairs from the summary and then using off-the-shelf QA models to check the consistency of answers extracted from the document. QuestEval also evaluates four dimensions beyond faithfulness and uses a weighter to identify important questions. QAFactEval (Fabbri et al., 2022) analyzes and combines the best components of the QA-based metric pipeline for improved performance, including answer selection, question generation, question answering, answer overlap evaluation, and question filtering. MQAG (Manakul et al., 2023a) computes the statistical distance of answers over automatically generated multiple-choice questions, allowing for a more abstract understanding of information. Counterfactual estimation, such as CoCo (Xie et al., 2021), measures the causal effects from source documents to generated summaries. It can also be generalized to multi-level granularity and other tasks. Other than QA-QG pipeline, FactGraph (Ribeiro et al., 2022) compares the generated summary with the input document through semantic graphs, thereby enhancing discourse comprehensibility. FactKB (Feng et al., 2023) introduces external triplet knowledge to augment the training data, and assess the consistency of news summarization. ExtEval (Zhang et al., 2023b) reveals that extractive summarization does not guarantee faithfulness, identifying five broad categories of unfaithfulness issues that the entailment model cannot fully address.
@@
For Dialog Generation, $\mathbf{Q^{2}}$ (Honovich et al., 2021) is a evaluation method for knowledge-grounded generative models. It incorporates QG, QA, and NLI models to measure the model’s output to reflect its consistency with its grounding source knowledge, and measures the answer span similarity using an NLI model. FactPush (Steen et al., 2023) provides a simple data augmentation method by adding phrases to sentences, which can improve the NLI model performance.
@@
For multiple tasks, AlignScore (Zha et al., 2023) can give a unified function for 7 tasks consistency evaluation. It breaks the context into chunks and the claim into sentences to calculate the alignment score. WeCheck (Wu et al., 2023) provides a weak supervision learning paradigm based on previous metrics and uses a mixture NLI datasets to warm up and do a noise-aware fine-tuning to get the target metric model.
@@
Compared to reference-based methods, these approaches have greater generality. It is noticeable that QG-QA models take a large portion. Because, without references, it is challenging for NLI models to get accurate extraction of evidence from the source text. Essentially, the QG-QA framework is a flexible approach that extracts and compares information from generated output and source input. However, the QG and QA steps may introduce considerable noise, which can interfere with the results. When comparing, some methods also incorporate entailment models for judgment. It is showed that entailment and QA-based metrics offer complementary signals that the combination can achieve better performance (Fabbri et al., 2022).